Suppose the resistance of the coil in the previous problem is 25 Ω. Assume that the coil moves with uniform velocity during its removal and restoration. Find the thermal energy developed in the coil during (a) its removal, (b) its restoration and (c) its motion.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. R = 25 Ω (a) e = 50V , T = 0.25 s I = e/R = 2A, H = i^2 RT = 4 x 25 x 0.25 = 25 J (b) e = 50 V, T = 0.25 s I = e/R = 2A, H = i^2 RT = 25 J (c) Since energy is a scalar quantity Net thermal energy developed = 25 J + 25 J = 50 J.